# Hierarchical modeling of gravitational-wave populations for disentangling environmental and modified-gravity effects

### v1

@@ I Introduction
Since the first detection of gravitational waves (GWs) in 2015 Abbott et al. (2016), the LIGO-Virgo-KAGRA (LVK) GW Transient Catalog (GWTC) has grown to include more than 200 transient sources Abbott et al. (2019a, 2021a, 2024, 2023); LIG (2025a) in the $\sim 100$ Hz band. The GWTC has been used to set tight constraints on the properties of merging compact binary populations in the universe LIG (2025b) and to test general relativity (GR) Abbott et al. (2019b, 2021b, 2021c). The upcoming network of third-generation ground-based detectors, including the Einstein Telescope Hild et al. (2011) and Cosmic Explorer Abbott et al. (2017); Reitze et al. (2019), will further improve these constraints and extend population studies to heavier binaries in the $\sim 1-10$ Hz band Singh et al. (2022); Ng et al. (2021); Gair et al. (2011a). At even lower frequencies $\sim 1$ mHz, the upcoming Laser Interferometer Space Antenna (LISA) will probe the properties of supermassive black hole (MBH) populations Colpi et al. (2024); Gair et al. (2011b).
@@
A particularly interesting class of sources in the LISA band is provided by extreme-mass-ratio inspirals (EMRIs), where a compact object (CO) of mass $\mu\sim 10^{1-2}M_{\odot}$ completes $\sim 10^{5}$ highly relativistic orbits around an MBH of mass $M\sim 10^{5-7}M_{\odot}$. EMRIs are expected to constrain MBH properties such as mass and spin to sub-percent precision Babak et al. (2017), enabling tight constraints on the MBH mass function Gair et al. (2010). Furthermore, the large number of strong field cycles completed by the CO makes EMRIs unique probes of astrophysical environments and potential deviations from GR Barausse et al. (2007); Barausse and Rezzolla (2008); Kocsis et al. (2011); Barausse et al. (2014, 2016); Arun et al. (2022); Cole et al. (2023); Duque et al. (2025a); Speri et al. (2023, 2024a); Copparoni et al. (2025). However, the sensitivity of EMRIs to such “beyond-vacuum-GR” effects may hinder their precision astronomy prospects. These effects are secular and perturbative in nature, often manifesting as simple power-law-like corrections to the EMRI’s vacuum-GR dynamics and introducing correlations across the parameter space. If these effects are present in the signal but are not included in the analysis, they can bias the source’s inferred parameters and potentially lead to, e.g., false claims of deviations from GR Speri et al. (2023); Kejriwal et al. (2023); Cárdenas-Avendaño and Sopuerta (2024).
@@
Previous studies have employed hierarchical models to search for evidence of modified gravity effects in GW populations (see, e.g., Li et al. (2012); Isi et al. (2019); Abbott et al. (2019b, 2021b, 2021c); Liu et al. (2024)), but do not consider multiple contending effects. Recently, Yuan et al. Yuan et al. (2025); Yuan (2025) proposed a statistic $F$ to distinguish modified gravity effects (with a constant universal amplitude) from environmental effects when they are present in the signal one at a time, but do not address the distinguishability of multiple concurrent effects in the system.
@@ II.2 Hypothesis testing
In the following, we set $r_{\alpha}=1$, effectively treating $\alpha(\vec{\lambda}^{\prime}|\mathcal{H}_{a})$ as a constant. As argued more formally in Appendix A, this assumption holds when $\vec{\lambda}^{\prime}$ is a set of beyond-vacuum-GR effect hyperparameters, e.g., the intrinsic strength of time-varying $G$. This is because such effects only perturbatively modify the detectability of a source in the population. Setting $r_{\alpha}=1$ mitigates some computational expense in the calculation of the Bayes factor, as the selection function involves an integral over source parameters over the full space. The integrand includes the detection probability $p_{\rm det}(\vec{\theta}|\mathcal{H})$, which is typically a numerical function in the GW context and hence expensive to compute. If $p_{\rm det}(\vec{\theta}|\mathcal{H})$ can be cheaply calculated, this factor may be restored feasibly in future studies (see, e.g., Chapman-Bird et al. (2023)).
@@ III.1 Source model
Here $p$ is the instantaneous semi-latus rectum of the CO’s orbit and $M$ is the MBH mass; $A_{\ell}$ is the effective strength of the migration torque effect, with typical values $\sim 10^{-5}$ Barausse et al. (2014); $n_{\ell}$ is the post-Newtonian (PN) order at which the local effect enters the evolution, and is $\approx 8$ for migration torques in $\alpha$-disks Barausse et al. (2014); $A_{g}$ is the strength of the time-varying $G$ effect, with current best constraints $\lesssim 10^{-12}~[\rm yr^{-1}]$ Chamberlain and Yunes (2017), and $n_{g}$ is its PN order in the evolution. For time-varying $G$, $n_{g}=4$, denoting a 4-4 PN order correction Chamberlain and Yunes (2017); Speri et al. (2023). Note that the flux balance equation reduces to vacuum-GR as $A_{\ell},A_{g}\to 0$. We emphasize that such parametric constructions are also available for other beyond-vacuum-GR effects in the literature Kocsis et al. (2011); Barausse et al. (2014, 2016); Arun et al. (2022); Cole et al. (2023); Duque et al. (2025a); Speri et al. (2024a); Kejriwal et al. (2023).
In addition to those described above, the EMRI model has the following parameters: $a$, the dimensionless spin of the black hole; $\mu$, the mass of the CO; $d_{L}$, the luminosity distance of the source, or equivalently the redshift $z\equiv z(d_{L})$; $(\theta_{S},\phi_{S})$, the sky localization angles; $(\theta_{K},\phi_{K})$, the spin orientation angles; and $\Phi_{0}$, the initial azimuthal orbital phase.
@@ III.2 Population model
where $\dot{G}~[\rm yr^{-1}]$ is the value of time-varying $G$ at the population level. Hence, the global effect hyperparameter $\vec{\lambda}_{g}=\{\dot{G}\}$. We stress that this distinction between the inferred source parameter $A_{g}$ and the hyperparameter $\dot{G}$ is important in the data analysis context, since the individual source likelihood only estimates the MLE for $A_{g}$, the value of which may be biased about the true value of $\dot{G}$ in the population depending on the assumed hypothesis, the detector-noise realization, etc. A constraint on $\dot{G}$ is only obtained hierarchically from these “noisy” estimates of $A_{g}$.
@@ V Results
With the hierarchical framework developed in Sec. II, the description of the EMRI model in Sec. III, and the data analysis setup in Sec. IV, we now perform hypothesis tests of beyond-vacuum-GR effects in various example cases. In the first subsection, we study how well our framework can recover the true hypothesis under which a given EMRI population was generated, considering three different population sizes. Then, in the second subsection, we show the dependence of the Bayes factors on the true local and global effect hyperparameters $f^{*}$ and $\dot{G}^{*}$ in the population. This will allow us to comment on the robustness of our framework, particularly in disentangling these effects. All results presented below are produced using the Hierarchical package, publicly available on Github (https://github.com/perturber/Hierarchical/).
@@ V.1.1 Setup
Detector response and GW observables — For a given source with true parameters $\vec{\theta}^{*}$, we generate a time-domain GW waveform $\vec{h}(\vec{\theta}^{*}):=\vec{h}_{+}(\vec{\theta}^{*})-i\vec{h}_{\times}(\vec{\theta}^{*})$ in the solar system barycenter (SSB) frame. We use the adiabatic Kerr equatorial inspiral model FastKerrEccentricEquatorial of the FastEMRIWaveforms (FEW) package Chua et al. (2021); Katz et al. (2021); Speri et al. (2024b); Chapman-Bird et al. (2025). Additionally, we fix the detector’s observation window to 1 year, such that sources that plunge earlier are observed until merger. Otherwise, their waveforms are cut off at the 1-year mark.
@@
This constitutes the set of $i=1,\ldots,N_{\rm obs}$ observed sources in the population. For the $i^{\rm th}$ source, we calculate the FIM in the full parameter space, $\Gamma_{i}$, using the StableEMRIFisher (SEF) package Kejriwal et al. (tion). The corresponding MLE estimate, $\hat{\theta}_{i}$, is obtained assuming the linear-bias approximation, Eq. (23).
@@ V.1.2 Results
Figure 4: The log-Bayes factor $\log_{10}\mathcal{B}^{g}_{\ell}$ as a function of $f^{*}$ (top panel) and $\dot{G}^{*}$ (bottom panel) in $11$ different populations of size $N_{\rm pop}=500$. In the top panel, $\dot{G}^{*}=10^{-12}[\rm yr^{-1}]$ is fixed. In the bottom panel, we fix $f^{*}=0.5$ instead. In the bottom panel, the dashed horizontal line represents $\log_{10}\mathcal{B}^{g}_{\ell}=0.0$, i.e., for which the data does not favour either of the hypotheses.
We now calculate the Bayes factors $\mathcal{B}^{v}_{\ell}$ and $\mathcal{B}^{v}_{g}$ according to Eqs. (18) and (19) (and $\mathcal{B}^{g}_{\ell}=\mathcal{B}^{v}_{\ell}/\mathcal{B}^{v}_{g}$). We again draw $N=5000$ samples of the vacuum-GR, local effect, and global effect hyperparameters, $\vec{\lambda}_{v},\vec{\lambda}_{\ell},\vec{\lambda}_{g}$, respectively, this time from their full priors (Table 4), and weight them by the approximate hyperlikelihoods calculated in the appendices B.2B.3, and B.4 for the vacuum-GR, local effect, and global effect hypotheses, respectively. We plot the log Bayes factors for all four populations and the different population sizes in Fig. 3. For brevity, we notate, e.g., $\log_{10}\mathcal{B}^{v}_{\ell}$ as $v/\ell$ in this plot. We find that the correct hypothesis is always preferred in all of the populations. Notably, in $P_{g}$, the local effect hypothesis is also preferred over vacuum-GR even though the true population does not have the local effect. This is again explained by the systematic biases induced on the MLE estimates of $A_{\ell},n_{\ell}$ due to a non-null $\dot{G}$ in this population. The two hypotheses $\mathcal{H}_{\ell},\mathcal{H}_{g}$ can still be disentangled by calculating $\mathcal{B}^{g}_{\ell}$, found in this case to be extremely favorable for the (correct) global effect hypothesis. This is represented by the $g/\ell$” bars in the third panel. The effect of the size of the EMRI population is most evident in the calculation of $\mathcal{B}^{g}_{\ell}$ in the population $P_{\rm mix}$, where the preference for the global effect hypothesis over the local one flips when $N_{\rm pop}$ is small. This is because with fewer sources in the population, the measurement uncertainty on $\dot{G}$ is larger, and hence increasingly consistent with the null value.
@@ V.2 Robustness check
To more broadly comment on the robustness of the hierarchical modeling framework for decoupling the considered beyond-vacuum-GR effects, i.e., migration torques and time-varying $G$, we now study the variability of the Bayes factor $\mathcal{B}^{g}_{\ell}$ as a function of the population’s hyperparameters.
@@
We consider 11 populations of fixed size $N_{\rm pop}=500$. In the first check, we fix $\dot{G}^{*}=10^{-12}[\rm yr^{-1}]$ and vary $f^{*}$ linearly in the range $[0.0,1.0]$. The setup otherwise is equivalent to that in the previous subsection. As shown in the top panel of Fig. 4, $\log_{10}(\mathcal{B}^{g}_{\ell})$, which quantifies the data’s preference for the global effect hypothesis over the local, scales inversely with $f^{*}$. This behavior is expected: as $f^{*}$ increases in the population, the MLE estimate of the global effect’s source parameter $\hat{A}_{g}$ is more biased among sources, and the inferred uncertainty in the hyperparameter $\dot{G}$ grows. This makes the hyperposterior on $\dot{G}$ more consistent with the null value, and hence the preference of the data for the global effect drops. We also calculate the relative error $r_{\dot{G}}:=|\dot{G}^{*}-\langle\dot{G}\rangle|/\dot{G}^{*}$ in all 11 populations. For $f\geq 0.9$, we find that $r_{\dot{G}}\gtrsim 1.0$, i.e., the expectation value $\langle\dot{G}\rangle$ incurs significant biases with respect to the truth $\dot{G}^{*}$. This is due to the biases induced by the local effect across a larger fraction of the population.
@@
In the second check, we fix $f^{*}=0.5$ and vary $\dot{G}^{*}$ linearly in the range $10^{-13},10^{-12}]$ across the 11 populations. As shown in the bottom panel of Fig. 4, $\log_{10}\mathcal{B}^{g}_{\ell}$ scales proportionately with the global effect hyperparameter $\dot{G}^{*}$. At smaller values, $\dot{G}^{*}\lesssim 4\times 10^{-13}~[\rm yr^{-1}]$, $\log_{10}\mathcal{B}^{g}_{\ell}\lesssim 0.0$, i.e., the global effect hypothesis is no longer favored. This is because the data is less sensitive to the global effect if it is too weak, such that the measurement uncertainty in $\dot{G}$ grows. Effectively, this analysis suggests that only population realizations with intrinsic strength of the global effect $\dot{G}^{*}\gtrsim 4\times 10^{-13}~[\rm yr^{-1}]$ will be discernible in this framework. We highlight that, at the zero crossing point of $\log_{10}\mathcal{B}^{g}_{\ell}$, even though both migration torques and time-varying $G$ are present in the population, they effectively “cancel” each other out. In such realizations of the population, the hierarchical framework will struggle to disentangle the two effects without supplementary checks.

### v2

@@ I Introduction
Since the first detection of gravitational waves (GWs) in 2015 Abbott and others (2016), the LIGO-Virgo-KAGRA (LVK) GW Transient Catalog (GWTC) has grown to include more than 200 transient sources B. P. Abbott et al. (2019a); R. Abbott et al. (2021a); R. Abbott et al. (2024); R. Abbott et al. (2023); 44 in the $\sim 100$ Hz band. The GWTC has been used to set tight constraints on the properties of merging compact binary populations in the universe 43 and to test general relativity (GR) Abbott and others (2019b, 2021b, 2021c). The upcoming network of third-generation ground-based detectors, including the Einstein Telescope Hild and others (2011) and Cosmic Explorer Abbott and others (2017); Reitze and others (2019), will further improve these constraints and extend population studies to heavier binaries in the $\sim 1-10$ Hz band Singh et al. (2022); Ng et al. (2021); Gair et al. (2011a). At even lower frequencies $\sim 1$ mHz, the upcoming Laser Interferometer Space Antenna (LISA) will probe the properties of supermassive black hole (MBH) populations Colpi and others (2024); Gair et al. (2011b).
@@
A particularly interesting class of sources in the LISA band is provided by extreme-mass-ratio inspirals (EMRIs), where a compact object (CO) of mass $\mu\sim 10^{1-2}M_{\odot}$ completes $\sim 10^{5}$ highly relativistic orbits around an MBH of mass $M\sim 10^{5-7}M_{\odot}$. EMRIs are expected to constrain MBH properties such as mass and spin to sub-percent precision Babak et al. (2017), enabling tight constraints on the MBH mass function Gair et al. (2010). Furthermore, the large number of strong field cycles completed by the CO makes EMRIs unique probes of astrophysical environments and potential deviations from GR Barausse et al. (2007); Barausse and Rezzolla (2008); Kocsis et al. (2011); Barausse et al. (2014, 2016); Arun and others (2022); Cole et al. (2023); Duque et al. (2025a); Speri et al. (2023, 2024a); Copparoni et al. (2025). However, the sensitivity of EMRIs to such “beyond-vacuum-GR” effects may hinder their precision astronomy prospects. These effects are secular and perturbative in nature, often manifesting as simple power-law-like corrections to the EMRI’s vacuum-GR dynamics and introducing correlations across the parameter space. If these effects are present in the signal but are not included in the analysis, they can bias the source’s inferred parameters and potentially lead to, e.g., false claims of deviations from GR Speri et al. (2023); Kejriwal et al. (2023); Cárdenas-Avendaño and Sopuerta (2024).
@@
Previous studies have employed hierarchical models to search for evidence of modified gravity effects in GW populations (see, e.g., Li et al. (2012); Isi et al. (2019); Abbott and others (2019b, 2021b, 2021c); Liu et al. (2024)), but do not consider multiple contending effects. Recently, Yuan et al. Yuan et al. (2025); Yuan (2025) proposed a statistic $F$ to distinguish modified gravity effects (with a constant universal amplitude) from environmental effects when they are present in the signal one at a time, but do not address the distinguishability of multiple concurrent effects in the system.
@@ II.2 Hypothesis testing
In the following, we set $r_{\alpha}=1$, effectively treating $\alpha(\vec{\lambda}^{\prime}|\mathcal{H}_{a})$ as a constant. As argued more formally in Appendix A, this assumption holds when $\vec{\lambda}^{\prime}$ is a set of beyond-vacuum-GR effect hyperparameters, e.g., the intrinsic strength of time-varying $G$. This is because such effects only perturbatively modify the detectability of a source in the population. Setting $r_{\alpha}=1$ mitigates some computational expense in the calculation of the Bayes factor, as the selection function involves an integral over source parameters over the full space. The integrand includes the detection probability $p_{\rm det}(\vec{\theta}|\mathcal{H})$, which is typically a numerical function in the GW context and hence expensive to compute. If $p_{\rm det}(\vec{\theta}|\mathcal{H})$ can be cheaply calculated, this factor may be restored feasibly in future studies (see, e.g., Chapman-Bird et al. (2023); Singh et al. (2025)).
@@ III.1 Source model
Here $p$ is the instantaneous semi-latus rectum of the CO’s orbit and $M$ is the MBH mass; $A_{\ell}$ is the effective strength of the migration torque effect, with typical values $\sim 10^{-5}$ Barausse et al. (2014); $n_{\ell}$ is the post-Newtonian (PN) order at which the local effect enters the evolution, and is $\approx 8$ for migration torques in $\alpha$-disks Barausse et al. (2014); $A_{g}$ is the strength of the time-varying $G$ effect, with current best constraints $\lesssim 10^{-12}~[\rm yr^{-1}]$ Chamberlain and Yunes (2017), and $n_{g}$ is its PN order in the evolution. For time-varying $G$, $n_{g}=4$, denoting a 4-4 PN order correction Chamberlain and Yunes (2017); Speri et al. (2023). Note that the flux balance equation reduces to vacuum-GR as $A_{\ell},A_{g}\to 0$. We emphasize that such parametric constructions are available for other beyond-vacuum-GR effects in the literature Kocsis et al. (2011); Barausse et al. (2014, 2016); Arun and others (2022); Cole et al. (2023); Duque et al. (2025a); Speri et al. (2024a); Kejriwal et al. (2023).
In addition to those described above, several other parameters define the EMRI model: $a$, the dimensionless spin of the black hole; $\mu$, the mass of the CO; $d_{L}$, the luminosity distance of the source, or equivalently the redshift $z\equiv z(d_{L})$; $(\theta_{S},\phi_{S})$, the sky localization angles; $(\theta_{K},\phi_{K})$, the spin orientation angles; and $\Phi_{0}$, the initial azimuthal orbital phase.
@@ III.2 Population model
where $\dot{G}~[\rm yr^{-1}]$ is the value of time-varying $G$ at the population level. Hence, the global effect hyperparameter $\vec{\lambda}_{g}=\{\dot{G}\}$. With this definition, while the inferred source parameter $A_{g}$ and the hyperparameter $\dot{G}$ are equivalent in the signal model, we stress that they may assume different values in the analysis. For example, in the global effect hypothesis, the inferred MLE of $A_{g}$ will be systematically biased by the true value of the local effect (migration torque) parameters in a given source. Since the local effect’s value changes across sources, the MLE $\hat{A}_{g}$ also becomes a source-dependent quantity. On the other hand, the inferred best-fit value of the hyperparameter $\dot{G}$, which is constant for all sources by definition, is inferred by “averaging” over these biased, source-dependent estimates of $A_{g}$. This averaging corresponds to the product over individual source likelihoods in the definition of the hyperposterior (Eq. (4)).
@@ V Results
With the hierarchical framework developed in Sec. II, the description of the EMRI model in Sec. III, and the data analysis setup in Sec. IV, we now perform hypothesis tests of beyond-vacuum-GR effects in various example cases. In the first subsection, we study how well our framework can recover the true hypothesis under which a given EMRI population was generated, considering three different population sizes. Then, in the second subsection, we show the dependence of the Bayes factors on the true local and global effect hyperparameters $f^{*}$ and $\dot{G}^{*}$ in the population. This will allow us to comment on the robustness of our framework, particularly in disentangling these effects. Finally, we briefly comment on the distinguishability of other beyond-vacuum-GR effects within the hierarchical framework. All results presented below are produced using the Hierarchical package, publicly available on Github (https://github.com/perturber/Hierarchical/).
@@ V.1.1 Setup
Detector response and GW observables — For a given source with true parameters $\vec{\theta}^{*}$, we generate a time-domain GW waveform $\vec{h}(\vec{\theta}^{*}):=\vec{h}_{+}(\vec{\theta}^{*})-i\vec{h}_{\times}(\vec{\theta}^{*})$ in the solar system barycenter (SSB) frame. We use the adiabatic Kerr equatorial inspiral model FastKerrEccentricEquatorial of the FastEMRIWaveforms (FEW) package Chua et al. (2021); Katz et al. (2021); Speri et al. (2024b); Chapman-Bird and others (2025). Additionally, we fix the detector’s observation window to 1 year, such that sources that plunge earlier are observed until merger. Otherwise, their waveforms are cut off at the 1-year mark.
@@
This constitutes the set of $i=1,\ldots,N_{\rm obs}$ observed sources in the population. For the $i^{\rm th}$ source, we calculate the FIM in the full parameter space, $\Gamma_{i}$, using the StableEMRIFisher (SEF) package Kejriwal et al. (manuscript in preparation). The corresponding MLE estimate, $\hat{\theta}_{i}$, is obtained assuming the linear-bias approximation, Eq. (24).
@@ V.1.2 Results
We now calculate the Bayes factors $\mathcal{B}^{v}_{\ell}$ and $\mathcal{B}^{v}_{g}$ according to Eqs. (18) and (19) (and $\mathcal{B}^{g}_{\ell}=\mathcal{B}^{v}_{\ell}/\mathcal{B}^{v}_{g}$). We again draw $N=5000$ samples of the vacuum-GR, local effect, and global effect hyperparameters, $\vec{\lambda}_{v},\vec{\lambda}_{\ell},\vec{\lambda}_{g}$, respectively, this time from their full priors (Table 4), and weight them by the approximate hyperlikelihoods calculated in the appendices B.2B.3, and B.4 for the vacuum-GR, local effect, and global effect hypotheses, respectively. We plot the log Bayes factors for all four populations and the different population sizes in Fig. 3. For brevity, we notate, e.g., $\log_{10}\mathcal{B}^{v}_{\ell}$ as $v/\ell$ in this plot. We find that the correct hypothesis is always preferred in all of the populations. Notably, in $P_{g}$ with $N_{\rm pop}=100,1000$, the local effect hypothesis is also preferred over vacuum-GR even though the true population does not have the local effect [^4] The deviation from the trend in the $N_{\rm pop}=500$ case may be attributed to one or more stochastic processes in the setup, which include generating the population, sampling from the hyperpriors, etc.. This is again explained by the systematic biases induced on the MLE estimates of $A_{\ell},n_{\ell}$ due to a non-null $\dot{G}$ in this population. The two hypotheses $\mathcal{H}_{\ell},\mathcal{H}_{g}$ can still be disentangled by calculating $\mathcal{B}^{g}_{\ell}$, found in this case to be extremely favorable for the (correct) global effect hypothesis. This is represented by the $g/\ell$” bars in the third panel. The effect of the size of the EMRI population is most evident in the calculation of $\mathcal{B}^{g}_{\ell}$ in the population $P_{\rm mix}$, where the preference for the global effect hypothesis over the local one flips when $N_{\rm pop}$ is small. This is because with fewer sources in the population, the measurement uncertainty on $\dot{G}$ is larger, and hence increasingly consistent with the null value.
@@ V.2 Robustness check
To more broadly comment on the robustness of the hierarchical modeling framework for decoupling the considered beyond-vacuum-GR effects, i.e., migration torques and time-varying $G$, we now study the variability of the Bayes factor $\mathcal{B}^{g}_{\ell}$ as a function of the population’s hyperparameters and size.
@@
We consider 11 populations of fixed sizes $N_{\rm pop}=100,500$. In the first check, we fix $\dot{G}^{*}=10^{-12}[\rm yr^{-1}]$ and vary $f^{*}$ linearly in the range $[0.0,1.0]$. The setup otherwise is equivalent to that in the previous subsection. As shown in the top panels of Fig 4, the log-Bayes factor $\log_{10}\mathcal{B}^{g}_{\ell}$ scales inversely with $f^{*}$ for both population sizes, as expected. While the $N_{\rm pop}=500$ population always prefers the global effect hypothesis, in the case of $N_{\rm pop}=100$, the log-Bayes factor crosses zero at $f^{*}\approx 0.5$. This is because as $f^{*}$ increases in the population, the MLE estimate of the global effect’s source parameter $\hat{A}_{g}$ is more biased across sources, and the inferred uncertainty in the hyperparameter $\dot{G}$ grows. This makes the hyperposterior on $\dot{G}$ more consistent with the null value, and hence the preference of the data for the global effect drops. In addition, we calculate the relative error $r_{\dot{G}}:=|\dot{G}^{*}-\langle\dot{G}\rangle|/\dot{G}^{*}$ in all 11 populations. For $f\geq 0.8$, $r_{\dot{G}}\gtrsim 1.0$ for $N_{\rm pop}=500$, i.e., the expectation value $\langle\dot{G}\rangle$ incurs significant biases with respect to the truth $\dot{G}^{*}$. This is due to the biases induced by the local effect across a larger fraction of the population. In the $N_{\rm pop}=100$ case, the larger measurement uncertainties cover any systematic biases sufficiently well, such that there are no significant biases.
@@
In the second check, we fix $f^{*}=0.5$ and vary $\dot{G}^{*}$ linearly in the range $10^{-13},10^{-12}]$ across the 11 populations. In populations of both sizes, $\log_{10}\mathcal{B}^{g}_{\ell}$ scales proportionally to the global effect hyperparameter $\dot{G}^{*}$ (bottom panel, Fig. 4). The log-Bayes factor $\log_{10}\mathcal{B}^{g}_{\ell}\lesssim 0.0$ for $N_{\rm pop}=500$ (=100=100) when $\dot{G}^{*}\lesssim 4\times 10^{-13}~[\rm yr^{-1}]$ ($\lesssim 10^{-12}~[\rm yr^{-1}]$), i.e., the global effect hypothesis is no longer favored. This is because the data is less sensitive to the global effect if it is too weak, such that the measurement uncertainty in $\dot{G}$ grows. In addition, at the zero crossing point of $\log_{10}\mathcal{B}^{g}_{\ell}$, even though both migration torques and time-varying $G$ are present in the population, they effectively “cancel” each other out. In such realizations of the population, the hierarchical framework will struggle to disentangle the two effects without supplementary checks. Finally, in the case of $N_{\rm pop}=100$, we note that the log-Bayes factors are much closer to zero because of the larger uncertainty in measuring $\dot{G}$.
Figure 4: The log-Bayes factor $g/\ell:=\log_{10}\mathcal{B}^{g}_{\ell}$ as a function of $f^{*}$ (top panel) and $\dot{G}^{*}$ (bottom panel) in $11$ different populations of size $N_{\rm pop}=500$ (orange circles) and $N_{\rm pop}=100$ (green triangles). In the top panels, $\dot{G}^{*}=10^{-12}[\rm yr^{-1}]$ is fixed and $f^{*}\in[0.0,1.0]$. In the bottom panel, we fix $f^{*}=0.5$ instead and $\dot{G}^{*}\in[10^{-13},10^{-12}]$. The horizontal line in both panels represents $\log_{10}\mathcal{B}^{g}_{\ell}=0.0$, i.e., for which the data does not favour either hypothesis.
V.3 A comment on other local effects
As was shown in Kejriwal et al. (2023), migration torques and time-varying $G$ can be severely degenerate in single source analyses across the EMRI parameter space. While other disk effects, such as accretion onto the MBH and dynamical friction Kocsis et al. (2011), may also strongly correlate with time-varying $G$, it was shown in Barausse et al. (2014) that migration torques induce the largest correction to the orbital phase of the system. This is why migration torques are the main focus of this paper. Since each effect induces a distinct PN-order correction to the source’s evolution, its measurability within the hierarchical framework may differ from that of migration torques and time-varying $G$ that we consider here, and should be studied in the future.
As another example, here we comment on the distinguishability of accretion (local effect) from time-varying $G$ (global effect) in the hierarchical framework. While accretion is not the dominant disk effect, it may be more severely degenerate with time-varying $G$ since both effects induce a correction to the fluxes at -4 PN order Kocsis et al. (2011); Barausse et al. (2014). We model the angular momentum flux as
$\dot{L}=\dot{L}_{\rm GR}\left(1+A^{\prime}_{\ell}\left(\frac{p}{M}\right)^{n^{\prime}_{\ell}}+A_{g}\left(\frac{p}{M}\right)^{n_{g}}\right)$~ (20)
where $A^{\prime}_{\ell}:=\dot{M}/M\lesssim 10^{-8}~[\rm yr^{-1}]$ is the accretion rate Kocsis et al. (2011) and $n^{\prime}_{\ell}=4$ corresponds to the -4 PN-order correction [^5] We use the primed notation to distinguish accretion effect parameters from the migration torque effect parameters considered in the rest of the paper.. We consider a population of size $N_{\rm pop}=500$ sources, with $f^{*}=0.2$ and $\dot{G}^{*}=0.0$. $A^{\prime*}_{\ell},n^{\prime*}_{\ell}$ are drawn from Eq. (13), with $\vec{\mu}^{*}_{\ell}=(10^{-10},4)$ and $(\sigma^{*}_{A^{\prime}_{\ell}},\sigma^{*}_{n^{\prime}_{\ell}})=(10^{-11},0.5)$.
In the analysis, we consider broad priors on $A^{\prime}_{\ell}\in[0.0,10^{-8}],n^{\prime}_{\ell}\in[-10.0,10.0]$, and initially keep the same priors on $A_{g}$ as in Table 3. The priors on the hyperparameters $f$ and $\dot{G}$ are also kept the same. Calculating the log-Bayes factors for the data’s preference for the three hypotheses, we find $\log_{10}\mathcal{B}^{v}_{\ell}\approx-20.83$, $\log_{10}\mathcal{B}^{v}_{g}\approx 1.31$, and $\log_{10}\mathcal{B}^{g}_{\ell}\approx-22.25$, i.e., the correct (local effect) hypothesis is preferred over the other two (vacuum and global), even when the local and the global effects are of the same PN order.
As a final robustness check, we now extend the analysis prior on the global effect amplitude to $A_{g}\in[0.0,10^{-8}]$, i.e., equal to the prior on $A^{\prime}_{\ell}$. Given the mathematically identical form of the two effects (Eq. (20)), these overlapping priors make the local and global effect parameters completely degenerate at the single source level. Recalculating the log-Bayes factors, we find that the local effect hypothesis is still preferred over the vacuum and the global ones, with $\log_{10}\mathcal{B}^{v}_{\ell}\approx-20.98$, $\log_{10}\mathcal{B}^{v}_{g}\approx 1.61$, and $\log_{10}\mathcal{B}^{g}_{\ell}\approx-22.59$. Although a more comprehensive analysis is beyond the scope of this work, this clearly demonstrates the hierarchical framework’s robustness in disentangling the local and global effect hypotheses over a single-source analysis, where the two hypotheses would be indistinguishable.